1. Sunshine Inc. has two equally-sized divisions. Division A has a beta of 0.8 and Division B has a beta of 1.2. The company is 100% equity financed. The risk-free rate is 6% and the market risk premium is 5%. Sunshine assigns different hurdle rates to each division based on each divisions market risk. Which of the following statements is CORRECT? Sunshines composite WACC is 10%. b. Division B has a lower WACC than Division A. C. If the same WACC is used for each division, the firm would select too many Division A projects and reject too many Division B projects. d. If the same WACC is used for each division, the firm would select too many Division B projects and reject too many Division A projects.
Answer #1

If the same WACC is used for each division, the firm would select too many Division B projects and reject too many Division A projects. (which is Option D)

_____

Explanation:

Option A is incorrect. The explanation is provided as below:

To calculate the composite WACC, we need to arrive at the beta for the company as a whole. The beta for the company is determined as below:

Composite Beta = Investment Percentage in Division A*Beta of Division A + Investment Percentage in Division B*Beta of Division B = 50%*.8 + 50%*1.2 = 1

Now, we can calculate the value of composite WACC as below:

WACC (Required Rate of Return) = Risk Free Rate + Composite Beta*Market Risk Premium

Substituting values in the above formula, we get,

WACC (Required Rate of Return) = 6% + 1*5% = 11%

____

Option B is incorrect. It is because the WACC of division B is higher than division A calculated as below:

WACC (Division A) = Risk Free Rate + Beta of Division A*Market Risk Premium = 6% + .8*5% = 10%

WACC (Division B) = Risk Free Rate + Beta of Division A*Market Risk Premium = 6% + 1.2*5% = 12%

Therefore, WACC of Division B (which is 12%) is greater than that of Division A (which is 10%).

____

Option C is incorrect. It is because projects offering returns which are less than composite WACC of 11% but greater than the WACC of Division A (10%) would get rejected resulting in acceptance of more risky projects (associated with Division B).

____

Option D is correct. It is because projects offering returns which are more than composite WACC of 11% but less than the WACC of Division B (12%) would get accepted resulting in rejection of less risky projects (associated with Division A).